A 90 day Commercial Bill with a Face Value of $50,000 was traded in the market 30 days after it was first issued. Current rates are: 30 days at 3.79% 60 days at 4.05% 90 days at 5.01% What price was the bill traded at ?
Answer #1

Face value = $50000

Trading date = 30 days after issuance

So days of bill now = 90-30= 60 days

60 days interest rate = 4.05%

Price of bill =Face value/(1+(rate *t/365))

=50000/(1+(4.05%*60/365))

=$49669.32477

So bond will trade at $49,669.32